OpenStudy (anonymous):

find a general solution to (D^4+4)y=0 please and promise i will guide u in a tour at elbaghl restaurant with chef president Ahmed Salmoun

OpenStudy (anonymous):

is that d to the fourth power?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

So, you're working on fourth order diff eqn?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

factor D^4+4 first,

OpenStudy (anonymous):

Are you there @zeki ?

OpenStudy (anonymous):

yeah with u ... i am just so exhausted sorry

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

So, did u find the factors?

OpenStudy (anonymous):

No i feel like being against a nuclear reactor

OpenStudy (anonymous):

Okay, let me help: D^4+4=D^4-(2i)^2=(D^2+2i)(D^2-2i)

OpenStudy (anonymous):

ooooooooooooooooh ! finally i got it .. so many thanks :) that is what i missed

OpenStudy (anonymous):

No probs :)